OpenStudy (anonymous):

find the other five trig ratios of angle. sec(angle)= 7/2

OpenStudy (anonymous):

for sin of angle i got sqrt 53/7 cos= 2/7 tan=sqrt53/2 csc= 7/sqrt 53 sec=7/2 cot= 2/sqrt 53

OpenStudy (sweetburger):